Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night Report 2018-02-14

Observers: Jesse T. Palmerio, Ken W. Smith,
Support Team: David O'Neill, Owen McBrien, Peter Clark, Erkki Kankare
TAT:  Cristina Barbarino, Morgan Fraser
TO: Mauricio Martinez (day) Eduardo Matamoros (night)

Calibrations: EFOSC2 PESSTO_calib, PESSTO_bias
Sky flats: g band sky flats in the evening

Weather: Thin, some clouds, sky with humidity 55%, variable seeing (0.8"-1.5").

Comments: Some difficulty in setting up the SOFI observations. One of the CCD quadrants had a gain problem. Wrong telluric for one object.

UT         Object         setup          exp          airmass  seeing  note
-----------------------------------------------------------------------------------------
EFOSC
23:40:00   Sky_flats                                                   Only g band, which was the one missing
00:28:00   Std_L745a                                           1.1"
00:57:00   SN2017ifu      gr13 sl1       2700s        1.8      1.1"
02:05:00   SN2018rw       gr13 sl1       1800s        1.03     1.1"    variable seeing (up to 1.5")
02:38:00   SN2018rw       gr13 sl1       1800s        1.06     1.3"    variable seeing (down to 0.8")
03:18:00   switch SOFI
SOFI
03:39:00   SN2018oh       JHK            32/32/60s    1.5      0.9" 
04:00:00   SN2018oh       BG sl1         480s         1.55     0.9"    No signal detected
04:40:00   Hip50489       BG RG sl1      4.5/6s       1.02     1.0"    We executed the telluric standard for SN2018ec instead of SN2018oh. Apologies.
04:50:00   SN2018oh       BG sl1         1620s        1.8-2.5  0.9"    Longer exposure 
06:30:00   Std_LTT3218                                         0.9"    SOFI flux standard
06:45:00   SN2018ec       BG sl1         1620s        1.1      0.8"    Problem with one quadrant of the CCD
08:30:00   Hip50489       BG RG sl1      4.5/6s       1.3      1.0"    Telluric standard for SN2018ec with problem on quadrant
08:32:00   SN2018ec       JHK            400/360/720s 1.32     1.0"    Problem with one quadrant of the CCD
09:09:00   switch EFOSC                                                Switch to EFOSC for morning spectroscopic arcs


Comments